Find the present value of $27 at 6 months, $58 at 2 years and $42 at 5 years if d(4)-8%.
Answer #1

Present Value (PV) =

1. 27 at 6 months

PV = 27 * PVF(8%,1/2)

PV = 27 * 0.96225 = 25.98

2. 58 at 2 years

PV = 58 * PVF(8%,2)

PV = 58 * 0.85734 = 49.73

2. 42 at 5 years

PV = 42 * PVF(8%,5)

PV = 42 * 0.68058 = 28.58